# supplementary

### v1

@@ 4 CITS Reveals Statistically Causal Neural Circuitry in the Mouse Brain During Visual Tasks
Figure 5: Stimulus-Specific Causal Functional Connectivity and Community Structure in Mouse Brain. A. Representative Causal Functional Connectivity (CFC) graphs for each stimulus type. Edges shown are those consistently present in over 80% of trials within each stimulus category (natural scenes, static gratings, and Gabor patches). Edge thickness indicates connection weight, transparency indicates edge stability across trials, and nodes (brain regions) are colored according to anatomical groupings. B. Strongly connected components (SCCs) derived from the corresponding CFC graphs, representing maximal directed communities in which each neuron can be reached from every other neuron via directed paths.
@@
4.3 Graph Motifs and Modular Substructure.
We investigated the strongly connected components (SCCs) identified in the CFC graphs for each stimulus type (Fig. 5B). SCCs represent tightly interlinked groups of nodes where each node is reachable from every other within the component, and their presence may reflect functional modules or recurrent processing motifs. Identifying such clusters helps uncover how local circuit dynamics reorganize in response to different sensory conditions, potentially reflecting stimulus-specific subnetworks involved in integration or feedback.
To identify SCCs, we applied a standard depth-first search (DFS)-based algorithm for detecting strongly connected components in directed graphs, implemented using the networkx Python package. Each CFC graph was analyzed separately by stimulus condition to extract the number, size, and node composition of SCCs, providing insight into stimulus-driven modular reconfiguration. These components represent mesoscale modules: clusters of neurons with mutual, cyclic influences that can be interpreted as candidate computational motifs or localized processing hubs [40].
Natural scenes exhibited the largest and most distributed SCCs, often bridging visual, thalamic, and hippocampal areas. This suggests that salient, complex, or sudden stimuli recruit widespread and integrative neural circuits, potentially reflecting the coordination of sensory, contextual, and memory-related processing [35, 41]. In contrast, static gratings and Gabor patches produced smaller, largely unimodal SCCs localized within primary and secondary visual cortices, consistent with their simpler spatial structure and more limited engagement of broader cortical networks [34]. These graph-theoretic motifs underscore how stimulus type shapes not only pairwise connectivity but also higher-order, recurrent causal structure, offering a systems-level perspective on how the brain dynamically reorganizes its functional modules in response to environmental input [40].
Additionally, the modular communities identified in each CFC graph (Fig. 5B, right panels) further highlight stimulus-specific clustering of causal interactions. Community structure was more fragmented for simpler stimuli, while natural scenes evoked more integrated modules spanning multiple brain areas, consistent with richer stimulus-driven coordination.
@@
Figure 6: Region-to-Region Distribution of Stable Causal Connections Across Visual Stimuli. Anatomical distribution of statistically causal edges across stimulus conditions. Bar plots quantify the percentage of neuron pairs exhibiting directed connections between each source and target region, based on edges that passed the 80% stability threshold. Regions are grouped into hippocampus (CA1, CA3, SUB), thalamus (LGv), and visual cortex (VISal, VISp). Bar heights indicate the percentage relative to all possible directed neuron pairs in each region group.
To further examine the anatomical structure and stimulus specificity of the inferred interactions, we quantified the proportion of directed edges originating from and targeting each brain region across stimulus types (Fig. 6). Natural scenes elicited the most widespread connectivity across all regions, including strong cortico-hippocampal and hippocampo-cortical interactions. VISp and VISal acted as dominant sources, sending projections to LGv, CA1, and SUB, while CA1 and SUB also showed recurrent intra-region connectivity. Static gratings produced a more localized pattern, with most edges confined to intra-visual cortex links and limited hippocampal involvement. Gabor stimuli yielded sparse connectivity overall, with weak interactions observed primarily within hippocampal subfields.
@@
4.5 The Relationship Between CFC and Visual Tuning Similarities
Figure 7: Connected Neurons Exhibit Greater Similarity in Direction and Orientation Selectivity. Direction selectivity index (DSI, left) and orientation selectivity index (OSI, right) were compared between neuron pairs that were either connected or unconnected based on the CITS-inferred CFC graphs. Only edges that passed the 80% trial-consistency threshold were considered connected. Across both static gratings and natural scenes, connected neurons exhibited lower DSI and OSI differences than unconnected pairs. Percentile ranks were computed relative to a null distribution of differences from 1000 random permutations.
Understanding whether causally connected neurons share similar tuning properties can reveal how functional specificity is embedded in the network architecture. In sensory systems, neurons with similar feature selectivity often form local subnetworks, potentially facilitating efficient encoding and processing of stimulus features [43]. Thus, examining the relationship between CFC and visual tuning offers insight into the functional organization of inferred networks. To explore this, we assessed whether neuron pairs connected in the CFC graph exhibited greater tuning similarity than unconnected pairs. Specifically, we calculated the differences in Orientation Selectivity Index (OSI) for all connected neuron pairs under each stimulus condition. To assess significance, we compared these differences to a null distribution generated from 1,000 randomly selected unconnected neuron pairs matched for neural activity range. (Fig. 7). For static gratings, connected neurons showed an OSI difference at the 1.9th percentile of the null distribution. For natural scenes, the OSI difference ranked at 7.5%. These findings suggest that connected neurons exhibit greater similarity in orientation tuning than unconnected neurons, with this effect more pronounced under static grating stimuli. We repeated these analysis for Direction Selectivity Index (DSI), and found that connected neurons show a DSI difference at the 63.8th percentile for static gratings and the 33.9th percentile for natural scenes. These results indicate that the connectivity derived by natural scenes or static gratings is more strongly associated with orientation selectivity than direction selectivity. This observation is consistent with prior anatomical and functional studies showing a strong “like-to-like” connectivity rule based on orientation preference in mouse visual cortex [44, 45], but not for direction selectivity. This suggests that orientation tuning is more tightly embedded in the anatomical structure of local circuits. This is consistent with prior studies showing that orientation selectivity is closely aligned with specific patterns of excitatory synaptic connectivity in visual cortex [43, 46], whereas direction selectivity is believed to arise more from circuit-level inhibition and feedforward processing than from long-range anatomical structure.
Overall, our findings suggest that CITS-inferred causal links capture not only directed statistical dependencies but also biologically relevant structure in functional tuning. While traditional functional connectivity approaches may reflect co-activation without directionality or selectivity [38], the alignment between CFC and tuning similarity points to a more structured, feature-sensitive organization. Taken together, our findings demonstrate that CITS uncovers putative directed neural circuits that are statistically robust, stimulus-specific, and physiologically interpretable across both micro- and meso-scales.
@@ Discussion
In contrast to Granger causality, which assumes linearity and can overfit noisy data, or PC-based algorithms, which are not designed for time series, CITS provides a unified, nonparametric approach tailored for temporally structured systems. The algorithm also extends beyond the recently proposed TPC framework [12] by avoiding assumptions of conditional DAGs within fixed windows and enabling inference under broader Markovian structures [51]. Whereas previous methods often struggle with latent confounding or nonlinear dependencies, CITS remains robust and accurate under both conditions, as shown in synthetic benchmarks and biological recordings. Moreover, the correlation between CITS-inferred edges and tuning similarity, as captured by DSI and OSI, reinforces its capacity to uncover functionally coherent circuits rather than merely statistical associations (see Fig. 7). This tuning-aligned connectivity pattern is consistent with earlier theoretical work linking information flow to feature-selective circuit motifs [52, 53].
@@
Despite these strengths, CITS has several limitations. First, CITS is formulated under the assumption of stationarity and fixed causal structure. While this is reasonable for short recordings or tightly controlled tasks, many biological systems exhibit non-stationary dynamics, including learning, adaptation, or state changes. Applying CITS over sliding or trial-based windows allows for coarse-grained tracking of such dynamics, but formal extensions to model time-varying causal graphs remain a compelling direction for future work. Second, the method depends on the accuracy and power of the conditional independence tests used. While we leverage Pearson partial correlation and the Hilbert-Schmidt Independence Criterion in our applications, these tests may lose sensitivity in high-dimensional or low-sample regimes. Third, the Markov order $\tau$ must be specified in advance. An adaptive or data-driven approach for estimating or integrating over $\tau$ could further improve performance and generalizability. A further methodological challenge is distinguishing intrinsic causal interactions from correlations induced by external stimuli  [54, 55]. Future extensions of CITS could incorporate stimulus regressors or leverage decorrelated stimulus designs to separate extrinsic from intrinsic sources of coupling.
@@
It is also important to emphasize that statistical causality does not necessarily imply direct physical connectivity. Inferred links may reflect indirect influences, shared upstream drivers, or coordinated dynamics across populations rather than monosynaptic projections. Integrating CITS with anatomical or interventional data, such as optogenetic stimulation, lesion studies, or tract-tracing, would provide additional validation and refinement. While we demonstrate the method in neural recordings, CITS is broadly applicable to any multivariate time series, including domains such as economics, climate systems, and behavioral ecology.
@@
Looking forward, several extensions and applications of CITS could deepen its utility. From a methodological perspective, incorporating latent variable models or dimensionality reduction may help scale to larger populations and reduce confounding. Introducing a Bayesian or regularized framework could improve performance in noisy or undersampled conditions. Applying CITS to multimodal datasets, such as joint electrophysiology and imaging, may enable cross-scale causal analysis. CITS also offers a framework for studying dynamic brain-wide interactions across a variety of cognitive and clinical contexts. For example, applying CITS to recordings from models of chronic pain, depression, or neuropsychiatric disorders could help uncover disruptions in top-down control circuits or altered causal flow across prefrontal, limbic, and sensory systems. This approach may offer a principled way to dissect circuit-level mechanisms underlying affective or cognitive dysfunction, potentially revealing new intervention targets grounded in statistically causal network architecture. Together, these directions position CITS as a foundational tool for discovering interpretable and robust causal structure in complex temporal systems.
@@ 5.2 Evaluation Metrics
[…]
@@ Acknowledgments
RA would like to acknowledge support from the National Institute of Mental Health of the National Institutes of Health under award number RF1MH128672 and the Sandler Program for Breakthrough Biomedical Research, which is partially funded by the Sandler Foundation.
@@
Lines 2-9 consider each possible edge $X_{u,s}\rightarrow X_{v,2\tau+1}$ and deletes the edge if for some $S\subseteq\{(d,r):d=1,\ldots,p;r=1,\ldots,2\tau+1\}$, $X_{u,s}\perp\!\!\!\!\perp X_{v,2\tau+1}|\bm{X}_{S}$, where $\bm{X}_{S}=\{X_{d,r}:(d,r)\in S\}$. This is justified by the implication (3$\implies$ (1) in Lemma 1 concluding that $X_{u,s}\not\in\text{pa}(v,2\tau+1)$.
@@
[…]
@@
using that the cardinality of the set |$|\{u,v,s,S\subseteq\mathbb{S}\}|=p^{2}\tau 2^{2p\tau-2}$. Then, for any $\gamma>0$, we have:
@@
$\sup_{u,v,s,S\subseteq\mathbb{S}}P(E_{u,v,s|S}^{II})\\ =\sup_{u,v,s,S\subset\mathbb{S}}P(|\hat{\mu}_{n}(X_{u,s},X_{v,2\tau+1}|S)|\leq\gamma\leavevmode\nobreak\ ,\\ \hskip 72.26999pt\mu(X_{u,s},X_{v,2\tau+1}|S)\neq 0)\\ \leq\sup_{u,v,s,S\subset\mathbb{S}}P(|\hat{\mu}_{n}(X_{u,s},X_{v,2\tau+1}|S)\\ \hskip 72.26999pt-\mu(X_{u,s},X_{v,2\tau+1}|S)|\geq c/2))\\ =o(1)\hskip 144.54pt.$ (7)
@@
$\rho(Y_{i},Y_{j}|Y_{\bm{k}})=0\text{ if and only if }Y_{i}\perp\!\!\!\!\perp Y_{j}\leavevmode\nobreak\ |\leavevmode\nobreak\ \{Y_{r}\leavevmode\nobreak\ :\ r\in\bm{k}\}.$
@@
Also, let $z(Y_{i},Y_{j}|Y_{\bm{k}})=g(\rho(Y_{i},Y_{j}|Y_{\bm{k}}))$. Note that $z(Y_{i},Y_{j}|Y_{\bm{k}})=0\Leftrightarrow\rho(Y_{i},Y_{j}|Y_{\bm{k}})=0$, and hence, $z(Y_{i},Y_{j}|Y_{\bm{k}})=0\Leftrightarrow Y_{i}\perp\!\!\!\!\perp Y_{j}\leavevmode\nobreak\ |\leavevmode\nobreak\ \{Y_{r}\leavevmode\nobreak\ :\ r\in\bm{k}\}$. That is, $z(Y_{i},Y_{j}|Y_{\bm{k}})$ is also a measure of conditional dependence. Furthermore, $Z_{n}(Y_{i},Y_{j}|Y_{\bm{k}})$ is a consistent estimator of the conditional dependence measure $z(Y_{i},Y_{j}|Y_{\bm{k}})$ (See Lemma III.1). Hence, it can be used to construct a conditional dependence test of the form $Z_{n}(Y_{i},Y_{j}|Y_{\bm{k}})>\gamma$ for some fixed $\gamma>0$, for the CITS in the Gaussian regime.
@@
[…]
    where $\epsilon_{i,t},i=1,\ldots,4,t=1,2,\ldots,1000,\sim$ i.i.d. $N(0,\eta)$ with mean 0 and standard deviation $\eta$. In this and all subsequent examples, the parameter $\eta$ is assumed to vary between 0 to 3.53.53.5 in increments of 0.50.50.5 in our numerical experiments. The true Rolled graph $G_{R}$ for this model has the edges 13,23,341$\rightarrow$ 3,2$\rightarrow$ 3,3$\rightarrow$ 41 → 3 , 2 → 3 , 3 → 4.
@@
[…]
    where $\epsilon_{i,t},i=1,\ldots,4,t=1,2,\ldots,1000,\sim$ i.i.d. $N(0,\eta)$ with mean 0 and standard deviation $\eta$. The true Rolled graph $G_{R}$ for this model has the edges 12,13,24,341$\rightarrow$ 2,1$\rightarrow$ 3,2$\rightarrow$ 4,3$\rightarrow$ 41 → 2 , 1 → 3 , 2 → 4 , 3 → 4.
@@
[…]
    where $\epsilon_{i,t},i=1,\ldots,4,t=1,2,\ldots,1000,\sim$ i.i.d. and uniformly distributed on the interval $(0,\eta)$. The true Rolled graph $G_{R}$ for this model has the edges 13,23,341$\rightarrow$ 3,2$\rightarrow$ 3,3$\rightarrow$ 41 → 3 , 2 → 3 , 3 → 4.
@@
[…]
    where $\epsilon_{i,t},i=1,\ldots,4,t=1,2,\ldots,1000,\sim$ i.i.d. Uniform distribution on the interval $(0,\eta)$. The true Rolled graph $G_{R}$ for this model has the edges 12,13,24,341$\rightarrow$ 2,1$\rightarrow$ 3,2$\rightarrow$ 4,3$\rightarrow$ 41 → 2 , 1 → 3 , 2 → 4 , 3 → 4.
@@
    We consider a motif consisting of $m=4$ and $w_{13}=w_{23}=w_{34}=10$ and $w_{ij}=0$ otherwise. We also note that in Eq. (12), $X_{j,t}$ depends on its own past. Therefore, the true Rolled graph has the edges 13,23,34,11,22,33,441$\rightarrow$ 3,2$\rightarrow$ 3,3$\rightarrow$ 4,1$\rightarrow$ 1,2$\rightarrow$ 2,3$\rightarrow$ 3,4$\rightarrow$ 41 → 3 , 2 → 3 , 3 → 4 , 1 → 1 , 2 → 2 , 3 → 3 , 4 → 4. The time constant $\tau_{i}$ is set to 10 units for each $i$. We consider $\epsilon_{i,t}$ to be distributed as an independent Gaussian process with mean 111 and standard deviation $\eta$. The signals are sampled at a time gap of $g:=\exp(1)\approx 2.72$ units for a total duration of 100010001000 units. For simulation purposes, one may replace the continuous derivative on the left hand side of Eq. (12) is replaced by first order differences at a gap of $g$.

### v2

@@ 4.2 Neural Signal Dependencies within Inferred CFC Motifs
Figure 6: Visual Stimulus-Specific Stable Causal Functional Connectivity and its Region-to-Region Distribution in the Mouse Brain. A. Representative Causal Functional Connectivity (CFC) graphs for each stimulus type. Edges shown are those consistently present in over 80% of trials within each stimulus category (natural scenes, static gratings, and Gabor patches). Edge thickness indicates connection weight, transparency indicates edge stability across trials, and nodes (brain regions) are colored according to anatomical groupings. B. Anatomical distribution of statistically causal edges across stimulus conditions. Bar plots quantify the percentage of neuron pairs exhibiting directed connections between each source and target region, based on edges that passed the 80% stability threshold. Regions are grouped into hippocampus (CA1, CA3, SUB), thalamus (LGv), and visual cortex (VISal, VISp). Bar heights indicate the percentage relative to all possible directed neuron pairs in each region group.
@@
To further examine the anatomical structure and stimulus specificity of the inferred interactions, we quantified the proportion of directed edges originating from and targeting each brain region across stimulus types (Fig. 6.B). Natural scenes elicited the most widespread connectivity across all regions, including strong cortico-hippocampal and hippocampo-cortical interactions. VISp and VISal acted as dominant sources, sending projections to LGv, CA1, and SUB, while CA1 and SUB also showed recurrent intra-region connectivity. Static gratings produced a more localized pattern, with most edges confined to intra-visual cortex links and limited hippocampal involvement. Gabor stimuli yielded sparse connectivity overall, with weak interactions observed primarily within hippocampal subfields.
@@ Discussion
In contrast to Granger causality, which assumes linearity and can overfit noisy data, or PC-based algorithms, which are not designed for time series, CITS provides a unified, nonparametric approach tailored for temporally structured systems. The algorithm also extends beyond the recently proposed TPC framework [8] by avoiding assumptions of conditional DAGs within fixed windows and enabling inference under broader Markovian structures [6]. Whereas previous methods often struggle with latent confounding or nonlinear dependencies, CITS remains robust and accurate under both conditions (see Theorem 2.2 and Corollary 1).
@@
Despite these strengths, CITS has several limitations. Firstly, the mathematical guarantees of CITS have been proved under stationarity of the underlying time series and fixed causal structure, that can be applicable to short recordings at high resolution. Many biological systems exhibit non-stationary dynamics, including learning, adaptation, or state changes. Applying CITS over sliding or trial-based windows yields an evolving CFC over time, but formal extensions to model time-varying causal graphs remain a compelling direction for future work. Second, the method relies on the accuracy and power of the conditional independence tests used. While we established that Pearson’s partial correlation and the Hilbert-Schmidt Independence Criterion based conditional dependence test yields accurate inference asymptotically over time series samples, it remains to establish the sensitivity of these tests in the high dimension low sample size regime. Third, the Markov order $\tau$ must be specified in advance or estimated based on estimation of order of stationary time series. Future extensions of CITS could incorporate stimulus regressors or leverage decorrelated stimulus designs to separate extrinsic from intrinsic sources of coupling.
@@
It is also important to emphasize that statistical causality does not necessarily imply direct anatomical synaptic connectivity. Inferred links may reflect shared upstream drivers, or coordinated dynamics across populations rather than monosynaptic projections. Integrating CITS with anatomical or interventional data, such as optogenetic stimulation, lesion studies, or tract-tracing, would provide additional validation and refinement. Furthermore, while we demonstrate CITS in the context of neural recordings, the method is broadly applicable to any multivariate time series, including domains such as economics, climate systems, and behavioral ecology.
@@
CITS also offers a framework for studying dynamic brain-wide interactions across a variety of cognitive and clinical contexts. For example, applying CITS to recordings from models of chronic pain, depression, or neuropsychiatric disorders could help uncover disruptions in top-down control circuits or altered causal flow across prefrontal, limbic, and sensory systems. This approach may offer a principled way to dissect circuit-level mechanisms underlying affective or cognitive dysfunction, potentially revealing new intervention targets grounded in statistically causal network architecture. Together, these directions position CITS as a foundational tool for discovering interpretable and robust causal structure in complex temporal systems.
@@ 5.2 Evaluation Metrics
which measures the proportion of correctly identified missing edges out of all true missing edges. Note that IFPR is scaled such that $100\%$ corresponds to perfect recovery of missing edges (i.e., no false positives).
@@ Acknowledgments
We thank members of the Abbasi Lab for their insightful feedback on this work. This study was partially supported by the National Institute of Mental Health of the National Institutes of Health under award number RF1MH128672 (original model development and initial validations), the National Library of Medicine of the National Institutes of Health under award number R01LM014619 (neuronal motif analysis), and the Sandler Program for Breakthrough Biomedical Research, which is partially funded by the Sandler Foundation. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
@@
[…]
@@
$P(\text{an error occurs in the sample CITS algorithm})$
@@
using that the cardinality of the set $|\{u,v,s,S\subseteq\mathbb{S}\}|=p^{2}\tau 2^{2p\tau-2}$. Then, for any $\gamma>0$, we have:
@@
$\sup_{u,v,s,S\subseteq\mathbb{S}}P(E_{u,v,s|S}^{II})\\ =\sup_{u,v,s,S\subset\mathbb{S}}P(|\hat{\mu}_{n}(X_{u,s},X_{v,2\tau+1}|S)|\leq\gamma\penalty 10000\ ,\\ \hskip 72.26999pt\mu(X_{u,s},X_{v,2\tau+1}|S)\neq 0)\\ \leq\sup_{u,v,s,S\subset\mathbb{S}}P(|\hat{\mu}_{n}(X_{u,s},X_{v,2\tau+1}|S)\\ \hskip 72.26999pt-\mu(X_{u,s},X_{v,2\tau+1}|S)|\geq c/2))\\ =o(1)\hskip 144.54pt.$ (7)
@@
$\rho(Y_{i},Y_{j}|Y_{\bm{k}})=0\text{ if and only if }Y_{i}\perp\!\!\!\!\perp Y_{j}\penalty 10000\ |\penalty 10000\ \{Y_{r}\penalty 10000\ :\ r\in\bm{k}\}.$
@@
Also, let $z(Y_{i},Y_{j}|Y_{\bm{k}})=g(\rho(Y_{i},Y_{j}|Y_{\bm{k}}))$. Note that $z(Y_{i},Y_{j}|Y_{\bm{k}})=0\Leftrightarrow\rho(Y_{i},Y_{j}|Y_{\bm{k}})=0$, and hence, $z(Y_{i},Y_{j}|Y_{\bm{k}})=0\Leftrightarrow Y_{i}\perp\!\!\!\!\perp Y_{j}\penalty 10000\ |\penalty 10000\ \{Y_{r}\penalty 10000\ :\ r\in\bm{k}\}$. That is, $z(Y_{i},Y_{j}|Y_{\bm{k}})$ is also a measure of conditional dependence. Furthermore, $Z_{n}(Y_{i},Y_{j}|Y_{\bm{k}})$ is a consistent estimator of the conditional dependence measure $z(Y_{i},Y_{j}|Y_{\bm{k}})$ (See Lemma III.1). Hence, it can be used to construct a conditional dependence test of the form $Z_{n}(Y_{i},Y_{j}|Y_{\bm{k}})>\gamma$ for some fixed $\gamma>0$, for the CITS in the Gaussian regime.
@@
    $\hskip 72.26999pt(X_{1,t},X_{2,t},X_{3,t},X_{4,t})=(1+\epsilon_{1,t},\\ \penalty 10000\ -1+\epsilon_{2,t},\penalty 10000\ 2X_{1,t-1}-X_{2,t-1}+\epsilon_{3,t},\penalty 10000\ 2X_{3,t-1}+\epsilon_{4,t})$
    where $\epsilon_{i,t},i=1,\ldots,4,t=1,2,\ldots,1000,\sim$ i.i.d. $N(0,\eta)$ with mean 0 and standard deviation $\eta$. In this and all subsequent examples, the parameter $\eta$ is assumed to vary between 0 to $3.5$ in increments of $0.5$ in our numerical experiments. The true Rolled graph $G_{R}$ for this model has the edges $1\rightarrow 3,2\rightarrow 3,3\rightarrow 4$.
@@
    $(X_{1,t},X_{2,t},X_{3,t},X_{4,t})=(1+\epsilon_{1,t},\penalty 10000\ -1+2X_{1,t-1}+\epsilon_{2,t},\\ \penalty 10000\ 2X_{1,t-1}+\epsilon_{3,t},\penalty 10000\ X_{2,t-1}+X_{3,t-1}+\epsilon_{4,t})$
    where $\epsilon_{i,t},i=1,\ldots,4,t=1,2,\ldots,1000,\sim$ i.i.d. $N(0,\eta)$ with mean 0 and standard deviation $\eta$. The true Rolled graph $G_{R}$ for this model has the edges $1\rightarrow 2,1\rightarrow 3,2\rightarrow 4,3\rightarrow 4$.
@@
    $\hskip 72.26999pt(X_{1,t},X_{2,t},X_{3,t},X_{4,t})=(\epsilon_{1,t},\\ \penalty 10000\ \epsilon_{2,t},\penalty 10000\ 4\sin(X_{1,t-1})-3\sin(X_{2,t-1})+\epsilon_{3,t},\penalty 10000\ 3X_{3,t-1}+\epsilon_{4,t})$
    where $\epsilon_{i,t},i=1,\ldots,4,t=1,2,\ldots,1000,\sim$ i.i.d. and uniformly distributed on the interval $(0,\eta)$. The true Rolled graph $G_{R}$ for this model has the edges $1\rightarrow 3,2\rightarrow 3,3\rightarrow 4$.
@@
    $(X_{1,t},X_{2,t},X_{3,t},X_{4,t})=(\epsilon_{1,t},\penalty 10000\ 4X_{1,t-1}+\epsilon_{2,t},\\ \penalty 10000\ 3\sin(X_{1,t-1})+\epsilon_{3,t},\penalty 10000\ 8\log(|X_{2,t-1}|)+9\log(|X_{3,t-1}|)+\epsilon_{4,t})$
    where $\epsilon_{i,t},i=1,\ldots,4,t=1,2,\ldots,1000,\sim$ i.i.d. Uniform distribution on the interval $(0,\eta)$. The true Rolled graph $G_{R}$ for this model has the edges $1\rightarrow 2,1\rightarrow 3,2\rightarrow 4,3\rightarrow 4$.
@@
    We consider a motif consisting of $m=4$ and $w_{13}=w_{23}=w_{34}=10$ and $w_{ij}=0$ otherwise. We also note that in Eq. (12), $X_{j,t}$ depends on its own past. Therefore, the true Rolled graph has the edges $1\rightarrow 3,2\rightarrow 3,3\rightarrow 4,1\rightarrow 1,2\rightarrow 2,3\rightarrow 3,4\rightarrow 4$. The time constant $\tau_{i}$ is set to 10 units for each $i$. We consider $\epsilon_{i,t}$ to be distributed as an independent Gaussian process with mean $1$ and standard deviation $\eta$. The signals are sampled at a time gap of $g:=\exp(1)\approx 2.72$ units for a total duration of $1000$ units. For simulation purposes, one may replace the continuous derivative on the left hand side of Eq. (12) is replaced by first order differences at a gap of $g$.